OpenStudy (anonymous):

|x - 1| = 5x + 10 I know the answers are x = -11/4 or x = -3/2 but how do i find out if one is extraneous or not?

OpenStudy (phi):

try each solution in the original equation and see if it works for example, with x= -11/4 |-11/4 - 1| =? 5* -11/4 +10 |-11/4 - 4/4| = ? -11/4 +40/4 |-15/4| = ? 29/4 15/4 = ? 29/4

OpenStudy (anonymous):

Ok thanks. Real quick, what happens to the 5 in step 2?

OpenStudy (phi):

I lost it! (-55+40)/4 = -15/4 but on the left side of the = you have | -15/4|= + 15/4 so it does not work

OpenStudy (phi):

in other words x= -11/4 ends up with +15/4 = ? - 15/4 which is not true on the other hand |-3/2 - 1 | = ? 5(-3/2) +10 5/2 = ? -15/2 + 20/2 5/2 = ? 5/2 which is true

OpenStudy (anonymous):

Oh ok thank you!

OpenStudy (phi):

yw